# Estimating Model Performance Under Covariate Shift Without Labels

### v2

@@ 4.2 Experimental setup
First, we fit machine learning models whose performance will be estimated. We separate the first-year data (2015) in each fetched dataset and use it as a training period. We fit 5 commonly used binary classification algorithms on the training data: Logistic Regression, Deep Learning model [7], Random Forest [15], XGBoost [2], and LGBM [10] with default parameters. We use these models to predict labels on the remaining part of the dataset. We record both - binary predictions and probability estimates. We end up with 1250 dataset-model pairs, which we call evaluation cases.
The rest of the data for each case is further divided into two periods - reference (the year 2016) and production (2017, 2018). Reference data is where the performance estimation methods are fitted using the monitored model inputs, outputs, and targets. Production data is further split into data chunks of 2000 observations each, maintaining the order of the observations. The realized performance of the monitored model is calculated on each chunk based on labels and monitored model outputs. The performance of the monitored model on each data chunk is also estimated based on monitored model inputs and outputs. The realized performance and the estimated performance are compared - each production chunk is an evaluation point of the performance estimation methods. After applying filtering that removed evaluation cases with less than 6000 observations (3 chunks) in the reference set and cases where the performance of the monitored model on the reference data was worse than random (that is with auroc lower than 0.5) we ended up with 959 evaluation cases. These contain 36 557 production data chunks (evaluation points) for each evaluated method. The metrics estimated are accuracy, auroc, and f1. Figure 1 shows an example of an auroc estimation result.
Figure 1: Estimation of auroc for ACSIncome problem data from the state of California and LGBM as the monitored model. The black line is the realized auroc of the monitored model for each chunk. The red line is the auroc estimated with $\mathrm{PAPE}$. The brown dashed line is the TEST SET performance.
We evaluate the performance estimation methods listed in 4.3 on all the production chunks in all the appropriate 959 evaluation cases [^1] We used a single 11th Gen Intel i7-11800H 2.30GHz machine, computation took over 120 hours.. The evaluation metrics are described, and the results are shown in 4.4.
4.3 Benchmarks
Below, we describe the benchmarks against which we evaluate $\mathrm{PAPE}$ and their implementation details. The code can be found in the paper’s GitHub repository [^2] https://github.com/pape-research/pape_r . We omit MANDOLINE from [1] as it requires user input on the nature of covariate shift.
@@ 4.3 Benchmarks
For each evaluation data chunk from production data, the performance estimated by this benchmark equals the performance calculated on reference data (and typically test set is chosen as the reference set). It is the main benchmark, representing the initial assumption under which the model’s performance on the production data is constant and equal to the performance calculated on test set.
@@
With ATC, we take the raw scores provided by the monitored model $f(x)$ and apply the maximum confidence function to them, denoting it as $\mathrm{MC}$:
$\mathrm{MC}\left(f(x)\right)=\begin{cases}f(x),&f(x)\geq 0.5\\ 1-f(x),&\text{otherwise}\end{cases}$ (7)
Then we use it to learn a threshold on reference data such that the fraction of observations above the threshold is equal to the performance metric value calculated on reference data. When inferring, we apply MC to the evaluation chunk and calculate the fraction of observations above the learned threshold. The estimated metric is equal to the calculated fraction.
@@
DoC was proposed to estimate image classifier accuracy [8]. The method assumes that the performance change is proportional to the change in mean of maximum confidence (7). To learn this relationship we fit Linear Regression on the difference of confidence between the shifted and the reference data as input and the difference in performance between these two as the target. In order to get datasets with synthetic distribution shifts we perform multiple random re-samplings of the reference dataset.
@@
Reverse Testing methods train a reverse model on production data with the monitored model inputs as features and monitored model predictions as targets. This reverse model is then used to make predictions on reference inputs, and its performance is evaluated with the reference labels. The reverse model’s performance on reference data is a proxy for the monitored model’s performance on the production data. We modify the method by additionally fitting Linear Regression to relate the reverse model performance change with the monitored model performance (just like in DoC).
@@
For IW, we first estimate density ratios $\widehat{w}(x)_{ref\rightarrow prod}$ between reference data and the production data chunks with the DRE classifier - the same that we use for $\mathrm{PAPE}$. Then we use them as weights to calculate the weighted performance metric of interest using reference data.
@@
CBPE is a simpler version of $\mathrm{PAPE}$. The key difference is that CBPE does simple calibration on reference data (without reweighing to production distribution).
Algorithm 2 Confidence-based Performance Estimation (CBPE)
  1. 1.
    Fit a naive calibrator $c$ on reference data with inputs $f(x_{ref})$ and targets $y_{ref}$
  2. 2.
    Use $c$ to calibrate $f$ scores on production data: $c\left(f(x_{prod})\right)$
  3. 3.
    Estimate performance metric $m$ on production data with equation 1.
4.4 Evaluation
Performance estimation is a regression problem. This leads to evaluation metrics such as mean absolute error (MAE) or root mean squared error (RMSE). However, the MAE/RMSE of all data chunks from multiple evaluation cases is meaningless. Large MAE/RMSE for an evaluation case where the performance of the model has a large variance might still be satisfactory. On the other hand, even small changes (in the absolute scale) in performance might be significant for cases where the model performance is very stable. To account for this, we scale absolute/squared errors by the standard error (SE) calculated for each evaluation case. We first measure SE as the standard deviation of the realized performance sampling distribution on the reference data. We do this by randomly sampling with repetitions 2000 observations (the size of the evaluation chunk) from the reference data and calculating the realized performance metric on each sample. We repeat this 500 times and calculate the standard deviation of the obtained per-sample metrics - the standard error (SE). Then, we divide absolute and squared errors by SE and squared SE, respectively. We call the SE-scaled metrics mean absolute standard error (MASTE) and root mean squared standard error (RMSSTE) (8).
@@ 4.4 Evaluation
Accuracy AUROC F1 MASTE RMSSTE MASTE RMSSTE MASTE RMSSTE TEST SET 1.68 2.96 1.48 2.33 2.54 8.24 RT-mod 2.36 5.51 1.89 4.33 2.05 4.71 ATC 1.58 2.73 1.91 3.49 3.02 9.23 DOC 1.18 1.92 1.39 2.78 2.58 8.58 CBPE 1.13 1.87 1.09 1.68 1.05 2.17 IW 1.05 1.42 1.07 1.56 1.06 1.70 PAPE 0.99 1.31 1.00 1.45 0.91 1.35
Table 1: MASTE and RMSSTE of the evaluated performance estimation methods for each estimated metric.
$\mathrm{PAPE}$ together with CBPE and IW, shows strong improvement over the TEST SET baseline. $\mathrm{PAPE}$ significantly outperforms the second best method for each estimation and evaluation metric (paired Wilcoxon signed-rank test $p=0.0$). $\mathrm{PAPE}$ result of MASTE equal to 0.99 for accuracy estimation means that the estimation is, on average, less than 1 SE away from the realized performance. For intuition, the MASTE of $\mathrm{PAPE}$ for the evaluation case shown in Figure 1 is equal 1.07. The TEST SET performance MASTE for accuracy is 1.68, indicating that accuracy changes significantly in the production chunks. If the performance was stable with only random normal fluctuations, the MASTE of the TEST SET performance would be around 0.8333That is because average absolute deviation for normal distribution is $\sqrt{\frac{2}{\pi}}\approx 0.8$ of its standard deviation.. The changes for f1 are much stronger, as shown by the increased MASTE of TEST SET - 2.49. This explains why performance estimation methods provide the strongest improvement compared to the TEST SET performance for this metric. Additionally, we have calculated MASTE and RMMSTE for each monitored model type - $\mathrm{PAPE}$ consistently outperformed IW for each of the monitored models and each estimation and evaluation metric.
We also analyze performance estimation algorithm’s accuracy across different levels of realized performance changes. For relatively small realized performance changes (within +/- 2SE), the TEST SET performance baseline is accurate. Other methods should match or exceed this baseline, particularly when realized performance changes are significant. We sorted the data by performance change magnitude and calculated rolling MASTE for 2SE-wide intervals. The results are shown in Figure 2.
Figure 2: Quality of performance estimation [MASTE of estimated metric] vs. absolute performance change [SE]. The x axis indicates the center of the data bucket - for example, value 1 indicates a bucket that contains data chunks for which the absolute performance change was in between 0 - 2 SE. The left y axis shows MASTE of the evaluated method for the data bucket. The right y axis shows the number of data chunks in each bucket on a logarithmic scale.
For all the estimated metrics, $\mathrm{PAPE}$ shows the best quality of estimations in the nearly whole evaluated region. It gets significantly better than other methods for large true performance changes. IW and CBPE show comparable performance, with IW being more accurate in most buckets.
@@ 4.5 Sample size effect
The experiments described in Section 4 were run for arbitrarily chosen sizes of the evaluation data chunk (2000 observations), so we have also investigated the quality of performance estimates for different chunk sizes. We split production data into chunks of the following sizes: 100, 200, 500, 1000, 2000, and 5000. The step, that is the difference between the beginning of each chunk, was set to 1000 observations for each chunk size. Due to computational complexity, we ran the experiment only for one evaluation case. We selected the biggest data set (California), the prediction task for which performance changes significantly (ACSEmployment), the default choice algorithm for the classification task on tabular data - LGBM, and the default metric - rocauc. Since the results come from a single evaluation case, we used a typical regression metric - mean absolute error (MAE). The results in Figure (3) show that all methods give less accurate auroc estimates for smaller chunk sizes. This is expected as the random noise effects are more significant for small samples. For the evaluated case, this has the strongest impact on IW accuracy of estimation.
@@
Figure 3: Effect of sample size on mean absolute error of auroc estimation. Calculated for sample sizes of 100, 200, 500, 1000, 2000, and 5000, on data from California, for prediction task: employment.
5 Conclusion
We introduced $\mathrm{PAPE}$, an innovative approach for accurately estimating the performance of classification models when faced with covariate shift. Our method was rigorously tested using diverse real-world tasks and datasets drawn from US Census data. The results demonstrated that $\mathrm{PAPE}$ significantly outperforms existing methods across all metrics assessed.
@@ 5 Conclusion
In addition, we developed a comprehensive new evaluation framework for performance estimation methods. This framework encompasses dataset preparation, benchmark implementation, and the introduction of novel evaluation metrics that are essential for aggregating results from multiple datasets and model monitoring scenarios. We employed this framework to assess $\mathrm{PAPE}$ alongside other methods, analyzing over 900 model-dataset pairs and generating more than 36,000 data chunks for thorough evaluation.
@@
Looking ahead, our future work will focus on quantifying the errors and uncertainties associated with these estimates and developing estimation intervals. A key source of error in our current methodology is the neglect of calibration errors, as indicated by $\mathrm{PAPE}$’s use of equation 1. To address this, we plan to implement calibration of $f$ within a cross-validation loop and estimate the conditional target probability $P(Y=1|s)$ for calibrated scores as a probability distribution, specifically using a Beta distribution. This will allow us to sample from this distribution for each $s$ and generate, using equation 1, a distribution of estimated performance metrics.

### v3

@@ 4.2 Experimental setup
First, we fit machine learning models whose performance will be estimated. We separate the first-year data (2015) in each fetched dataset and use it as a training period. We fit 5 commonly used binary classification algorithms on the training data: Logistic Regression [16], neural network model [17], Random Forest [16], XGBoost [18], and LGBM [19] with default parameters. We use these models to predict labels on the remaining part of the dataset. We record both - binary predictions and probability estimates. We end up with 1250 dataset-model pairs, which we call evaluation cases.
The rest of the data for each case is further divided into two periods - reference (the year 2016) and production (2017, 2018). Reference data is where the performance estimation methods are fitted using the monitored model inputs, outputs, and targets. Production data is further split into data chunks of 2000 observations each, maintaining the order of the observations. The realized performance of the monitored model is calculated on each chunk based on labels and monitored model outputs. The performance of the monitored model on each data chunk is also estimated based on monitored model inputs and outputs. The realized performance and the estimated performance are compared - each production chunk is an evaluation point of the performance estimation methods. After applying filtering that removed evaluation cases with less than 6000 observations (3 chunks) in the reference set and cases where the performance of the monitored model on the reference data was worse than random (with AUROC lower than 0.5 or F1 equal 0), we ended up with 959 evaluation cases. These contain 36 675 production data chunks (evaluation points) for each evaluated method. The metrics estimated are accuracy, AUROC, and F1. Figure 1 shows an example of an AUROC estimation result.
Figure 1: Estimation of AUROC for ACSIncome task data from California and LGBM as the monitored model. The black line is the realized AUROC of the monitored model. The red line and brown dashed line are estimates respectively from PAPE and the TEST SET performance benchmark.
We evaluate the performance estimation methods listed in 4.3 on all the production chunks in all the appropriate evaluation cases [^1] We ran the experiments on a single Intel i7-11800H 2.30GHz, 32GB RAM machine. Data preparation took about 48 hours (partially limited by internet bandwidth) while evaluating the methods took over 29 hours.. The evaluation metrics are described, and the results are shown in 4.4.
4.3 Evaluated methods
Below, we briefly describe the evaluated methods. We omit MANDOLINE from [7] as it requires user input on the nature of covariate shift.
PAPE
We implement PAPE following algorithm 1. In step 1 we estimate the weights $w$ with DRE as described in [20, chapter 2.7.5]. We use LGBM Classifier with default hyperparameters for this. In steps 2 and 3, we approximate weighted calibration under reference distribution and weights $w$ by fitting $w$-weighted regressor on reference data. We use LGBM Regressor with default hyperparameters.
@@ 4.3 Evaluated methods
For each evaluation data chunk from production data, the estimated performance equals the performance calculated on reference data (and typically, the test set is chosen as the reference set, hence the name). It represents the initial assumption under which the model’s performance on the production data is constant and equal to the performance calculated on the test set.
@@
With ATC, we take the raw scores provided by the monitored model $f(x)$ and apply the maximum confidence function to them, that is: $\mathrm{MC\bigl{(}f(x)\bigr{)}}=f(x)$ when $f(x)\geq 0.5$ and $1-f(x)$ otherwise. Then, we use it to learn a threshold on reference data such that the fraction of observations above the threshold equals the performance metric value calculated on reference data. When inferring, we apply MC to the evaluation chunk and calculate the fraction of observations above the learned threshold. The estimated metric is equal to the calculated fraction.
@@
DoC was proposed to estimate image classifier accuracy [10]. The method assumes that the performance change is proportional to the change in the mean of maximum confidence. To learn this relationship, we fit Linear Regression on the mean difference of confidence between the shifted and the reference data as inputs and the difference in performance between these two as the targets. In order to get datasets with synthetic distribution shifts we perform multiple random re-samplings of the reference dataset.
@@
Reverse Testing methods train a reverse model on production data with the monitored model inputs as features and monitored model predictions as targets. This reverse model is then used to make predictions on reference inputs, and its performance is evaluated with the reference labels. The reverse model’s performance on reference data is a proxy for the monitored model’s performance on the production data. We modified the method by finding a constant to ensure that the average performance estimate from the reverse model on reference data matches the realized performance. Without this modification the estimates returned by the method were always biased - either over or underestimating the performance metric.
@@
For IW, we first estimate density ratios $\widehat{w}(x)_{ref\rightarrow prod}$ between reference data and the production data chunks with the DRE classifier - the same we use for PAPE. Then, we use them as weights to calculate the weighted performance metric of interest using reference data.
@@
CBPE is a simpler version of PAPE. The difference is that CBPE calibrates on reference data without reweighing to production distribution, so it does not fully account for covariate shift impact.
4.4 Results
Performance estimation is a regression problem. This leads to evaluation metrics such as mean absolute error (MAE) or root mean squared error (RMSE). However, the MAE/RMSE of all data chunks from multiple evaluation cases is meaningless. Large MAE/RMSE for an evaluation case where the performance of the model has a large variance might still be satisfactory. On the other hand, even small changes (in the absolute scale) in performance might be significant for cases where the model performance is very stable. To account for this, we scale absolute/squared errors by the standard error (SE) calculated for each evaluation case and performance metric. We first measure SE as the standard deviation of the realized performance sampling distribution on the reference data. We do this by randomly sampling with repetitions 2000 observations (the size of the evaluation chunk) from the reference data and calculating the realized performance metric on each sample. We repeat this 500 times and calculate the standard deviation of the obtained per-sample metrics - the standard error. Then, we divide absolute and squared errors by SE and squared SE, respectively. We call the SE-scaled metrics mean absolute standard error (MASTE) and root mean squared standard error (RMSSTE) (7).
@@ 4.4 Results
Table 1: MASTE and RMSSTE of the evaluated methods for each estimated metric.
Accuracy AUROC F1 MASTE RMSSTE MASTE RMSSTE MASTE RMSSTE TEST SET 1.62 2.88 1.45 2.30 2.53 8.23 RT-mod 2.31 5.41 1.85 4.29 2.06 4.74 ATC 1.58 2.79 1.90 3.52 2.97 9.06 DOC 1.13 1.80 1.37 2.75 2.56 8.52 CBPE 1.08 1.75 1.07 1.68 1.03 2.12 IW 1.04 1.40 1.06 1.56 1.07 1.74 PAPE 0.97 1.28 0.99 1.45 0.90 1.34
PAPE, together with CBPE and IW, shows strong improvement over the TEST SET baseline. PAPE significantly outperforms the second-best method for each estimation and evaluation metric (paired t-test [21] p-value $=0.00$). PAPE result of MASTE equal to 0.97 for accuracy estimation means that the estimation is, on average, less than 1 SE away from the realized performance. For intuition, the MASTE of PAPE for the evaluation case shown in Figure 1 is equal 1.11. The TEST SET performance MASTE for accuracy is 1.62, indicating that accuracy changes significantly in the production chunks. If the performance was stable with only random normal fluctuations, the MASTE of the TEST SET performance would be around 0.8. That is because the average absolute deviation for normal distribution is $\sqrt{\frac{2}{\pi}}\approx 0.8$ of its standard deviation. The changes for F1 are much stronger, as shown by the increased MASTE of TEST SET - 2.53. This explains why performance estimation methods provide the strongest improvement compared to the TEST SET performance for this metric.
We also analyze the quality of estimations across different levels of realized performance changes. The TEST SET performance baseline is accurate for relatively small realized performance changes (within +/- 2SE). Other methods should match or exceed this baseline, particularly when realized performance changes are significant. We sorted the data by absolute performance change magnitude and calculated rolling MASTE for 2SE-wide intervals. The results are shown in Figure 2.
Figure 2: Quality of performance estimation [MASTE of estimated metric] vs. absolute performance change [SE]. The x axis indicates the center of the data bucket - for example, value 1 indicates a bucket that contains data chunks for which the absolute performance change was in between 0 - 2 SE. The left y axis shows MASTE of the evaluated method for the data bucket. The right y axis shows the number of data chunks in each bucket on a logarithmic scale. Methods not included in the plots (like RT-mod in plot A) performed significantly worse. Including them would change the scale of the plot and make comparison between best-performing solutions difficult.
For all the estimated metrics, PAPE estimation error is the lowest in nearly whole evaluated region. IW and CBPE show comparable accuracy.
@@ 4.5 Sample size effect
The experiments described above were run for arbitrarily chosen size of the evaluation data chunk (2000 observations), so we have also investigated the quality of performance estimates for different chunk sizes. We split production data into chunks of the following sizes: 100, 200, 500, 1000, 2000, and 5000. The step, that is the difference between the beginning of each chunk, was set to 1000 observations for each chunk size. Due to computational complexity, we ran the experiment only for one evaluation case  [^2] Computation took over 9 hours.. We selected the biggest data set (California), the prediction task for which performance changes significantly (ACSEmployment), the default choice algorithm for the classification task on tabular data - LGBM, and the default metric - AUROC. Since the results come from a single evaluation case, we used mean absolute error (MAE). The results in Figure (3) show that all methods give less accurate AUROC estimates for smaller chunk sizes. This is expected as the random noise effects are more significant for small samples. For the evaluated case, this has the strongest impact on IW accuracy of estimation.
@@
Figure 3: Effect of sample size on mean absolute error of AUROC estimation. Calculated for sample sizes of 100, 200, 500, 1000, 2000, and 5000, on data from California, prediction task ACSEmployment, client model: LGBM.
5 Conclusions
@@
We introduced PAPE, a novel approach for accurately estimating the performance of classification models without labels when faced with covariate shift. We developed an evaluation methodology for performance estimation methods. This methodology encompasses dataset preparation, benchmark implementation, and the introduction of new evaluation metrics that are essential for aggregating results from multiple datasets and model monitoring scenarios. We employed this evaluation methodology to assess PAPE alongside other methods, analyzing over 900 model-dataset pairs and generating more than 36,000 data chunks for thorough evaluation. The results show that PAPE provided more accurate performance estimates than other evaluated methods.
@@
Looking ahead, our future work will focus on quantifying the errors and uncertainties associated with these estimates and developing estimation intervals. A key source of error in our current methodology is the neglect of calibration errors, as indicated by PAPE’s use of equation 1. To address this, we plan to implement the calibration of $f$ within a cross-validation loop and estimate the conditional target probability $P(Y=1|s)$ for calibrated scores as a probability distribution, specifically using a Beta distribution. This will allow us to sample from this distribution for each $s$ and generate, using equation 1, a distribution of estimated performance metrics. Apart from that, we are actively looking for other sources of relevant datasets to extend our evaluation and better assess performance estimation methods.
Acknowledgments and Disclosure of Funding
The research described in this paper was done within NannyML - a venture-backed open-source software company focusing on post-deployment data science solutions. The company has received public R&D funding from Flanders Innovation & Entrepreneurship (VLAIO) - project number: HBC.2022.0846.
@@
Appendix A Appendix / Estimating metrics other than accuracy
Let us recall equation (1) of the PAPE algorithm that allows to estimate performance metric $m$ on production data based on monitored model predictions $g\bigl{(}f(x)\bigr{)}$ and calibrated probabilities $c\bigl{(}f(x)\bigr{)}$:
$\widehat{m}_{prod}=\mathrm{mean}_{x\sim prod}\left[c\bigl{(}f(x)\bigr{)}m\bigl{(}g(f(x)),1\bigr{)}+\bigl{(}1-c(f(x))\bigr{)}m\bigl{(}g(f(x)),0\bigr{)}\right]$
This equation can be used directly to estimate a performance metric that can be calculated as a mean of observation-level values, like accuracy. For metrics based on confusion matrix elements like F1, precision, recall and other, we first estimate these elements’ expected values. Generally function $g$ is any transformation that turns scores $f(x)$ into actionable predictions. For performance metrics based on confusion matrix $g$ needs to return binary predictions - one of which is to be considered positive and the other - negative. Typically $g\bigl{(}f(x)\bigr{)}$ returns 00 or $1$ and we consider these predictions, respectively, negative and positive. Assuming that is the case we estimate confusion matrix elements with:
$\widehat{tp}_{prod}$ $=\mathrm{mean}_{x\sim prod}\Bigl{(}c\bigl{(}f(x)\bigr{)}\cdot g\bigl{(}f(x)\bigr{)}\Bigr{)}$
$\widehat{fp}_{prod}$ $=\mathrm{mean}_{x\sim prod}\Bigl{(}\Bigl{(}1-c\bigl{(}f(x)\bigr{)}\Bigr{)}\cdot g\bigl{(}f(x)\bigr{)}\Bigr{)}$
$\widehat{tn}_{prod}$ $=\mathrm{mean}_{x\sim prod}\Bigl{(}c\bigl{(}f(x)\bigr{)}\cdot\Bigl{(}1-g\bigl{(}f(x)\bigr{)}\Bigr{)}\Bigr{)}$
$\widehat{fn}_{prod}$ $=\mathrm{mean}_{x\sim prod}\Bigl{(}\Bigl{(}1-c\bigl{(}f(x)\bigr{)}\Bigr{)}\cdot\Bigl{(}1-g\bigl{(}f(x)\bigr{)}\Bigr{)}\Bigr{)}$
Having these we can, for example, estimate recall with:
$\widehat{\mathrm{recall}}_{prod}\approx\frac{\widehat{tp}_{prod}}{\widehat{tp}_{prod}+\widehat{fn}_{prod}}$
To calculate (and estimate) AUROC we do not need model binary predictions $g\left(f(x)\right)$ instead we use raw scores $f(x)$ returned by the classifier. We define a new function $h_{t}$:
$h_{t}\bigl{(}f(x),t\bigr{)}=\begin{cases}1,&f(x)\geq t\\ 0,&f(x)<t\end{cases}$
Using this function we estimate confusion matrix elements for all distinct scores returned by the classifer, that is for all $t\in S$. Having those we estimate true positive rates (TPR) and false positive rates (FPR). Finally, we calculate the area under the curve created from estimated TPRs and FPRs to get AUROC.
Notice that the formula for recall gives only approximate result. The same applies to TPRs and FPRs obtained in the described way. That is because in general case the quotient of expected values of the random variables is not equal to the expected value of the quotient. In order to get an actual expectation of confusion-matrix-based metric we could perform sampling, create its distribution and calculate the expected value. We did not do this because it is computationally expensive, and the approximate formula gave accurate enough results for the metrics we experimented with (F1 and AUROC). In future research we aim to address this problem by providing efficient solution to calculate unbiased expectation of the performance metrics of interest.